OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (1, -4)with a slope of

OpenStudy (anonymous):

with a slope of 1/2

OpenStudy (anonymous):

The equation for this type of problem is (y-y1)=m(x-x1) with m as the slope. Plugging in you get: y--4=(1/2)(x-1) which simplifies to y+4=.5(x-1) and you can simplify it further if need be.